1.The waiting time, in hour, between successive speeders spotted by a radar units is a continuous random variable with cumulative distribution function

"f(x)={(1-e^(-8x), if x>0,0 otherwise)"

 derive the characteristic function of x and use it to find the mean of x


1
Expert's answer
2022-05-09T11:13:23-0400

The characteristic function has the form: "\varphi(t)=E[e^{itX}]". We receive the following expression: "\varphi(t)=\int_{-\infty}^{+\infty}e^{itx}p(x)dx", where "t\in{\mathbb{R}}". "p(x)" is the probability density function. "p(x)=f'(x)=8e^{-8x}", "x>0". For "x<0" the probability density function is . We get: "\varphi(t)=8\int_{0}^{+\infty}e^{itx-8x}dx=\frac{8}{it-8}e^{itx-8x}|_0^{+\infty}=-\frac{8}{it-8}". It is known that: "\frac{\partial\varphi(t)}{\partial t}|_{t=0}=iE[X]". We get: "\frac{\partial\varphi(t)}{\partial t}=\frac{8i}{(it-8)^2}". We receive: "E[X]=\frac18".

Answer: "E[X]=\frac18."